# Polyhedral Complex Derivation from Piecewise Trilinear Networks

### v1

@@ Abstract
Polyhedral Complex, Neural Radiance Fields, 3D Mesh
[…]
@@ 1 Introduction
Recent advancements in visualizing deep neural networks (Zhang et al., 2020; Lei & Jia, 2020; Lei et al., 2021; Berzins, 2023) contribute significantly to understanding the intricate structures that define these networks. This progress not only provides valuable insights into their expressivity, robustness, training methodologies, and distinctive geometry, but it also, by leveraging the inherent piecewise linearity in Continuous Piecewise Affine (CPWA) functions, e.g., ReLU neural networks, represents each region as a convex polyhedral, and the assembly of these polyhedral sets constructs a polyhedral complex that delineates the decision boundaries of neural networks (Grigsby & Lindsey, 2022).
@@
Building upon these breakthroughs, we explore the exciting prospect of analytically extracting a mesh representation from neural implicit surface networks (Wang et al., 2021; Yariv et al., 2023; Li et al., 2023). The extraction of a mesh not only offers a precise visualization and characterization of the network’s geometry but also does so efficiently through the utilization of vertices and faces, in contrast to sampling-based methods (Lorensen & Cline, 1987; Shen et al., 2021).
@@
These networks typically learn a signed distance function (SDF) by utilizing ReLU neural networks. However, recent approaches incorporate non-linear positional encoding techniques, such as trigonometric functions (Mildenhall et al., 2021) or trilinear interpolations using hashing (Müller et al., 2022) or tensor factorization (Chen et al., 2022), to mitigate issues like spectral bias (Tancik et al., 2020), ensuring fast convergence, and maintaining high-fidelity. Consequently, applying mesh extraction techniques based on CPWA functions becomes challenging.
Figure 1: The normal map and skeleton visualization are analytically generated from piecewise trilinear networks, comprising both HashGrid (Müller et al., 2022) and ReLU neural networks, that have been trained to learn a signed distance function with the eikonal loss for the Stanford bunny (Curless & Levoy, 1996).
@@
Focusing on the successful and widely-adopted trilinear interpolating methods, we present novel theoretical insights and a practical methodology for precise mesh extraction. Following a novel review of the edge subdivision method (Berzins, 2023), an efficient mesh extraction technique for CPWA functions, within the framework of tropical geometry (Section 3), we demonstrate that ReLU neural networks incorporating the trilinear module are piecewise trilinear networks (Definition 4.1). Then, we establish that the hypersurface within the trilinear region becomes a plane under the eikonal (Bruns, 1895) constraint (Theorem 4.5), a common practice during the training of SDF.
@@
Building upon this observation, we propose a method for approximating the determination of intersecting points among three hypersurfaces. Given the inherent curvature of these surfaces, achieving an exact solution is infeasible. In our approach, we utilize two hypersurfaces and a diagonal plane for a more feasible and precise approximation (Theorem 4.7). Additionally, we introduce a methodology for arranging vertices composed of faces to implicitly represent normal vectors. We experimentally confirm accuracy and simplicity using chamfer distance, efficiency, and angular distance, simultaneously exploring the relationship between the eikonal loss and the planarity of the hypersurface. In Figure 1, we present the analytically generated normal map and skeleton visualizations produced by our method using piecewise trilinear networks composed of HashGrid and ReLU neural networks. For a more detailed view and comprehensive explanations, please refer to Figure 7 in Appendix.
As we delve into this novel avenue, we not only aim to extend the existing visualization techniques to trilinear neural networks but also to explore potential new directions stemming from the applications of mesh extraction, e.g., baking for real-time rendering or geometric loss using extracted geometry (Berzins, 2023) in future work.
@@
    We present novel theoretical insights and a practical methodology for precise mesh extraction, employing piecewise trilinear networks.
@@
    We empirically validate correctness and parsimony through chamfer distance and efficiency, and angular distance, while examining the correlation between the eikonal loss and hypersurface planarity.
@@ 2 Related Work
A conventional black-box approach is to assess a given function at sampled points to derive the complex of the function, commonly implemented through marching cubes (Lorensen & Cline, 1987) or marching tetrahedra (Doi & Koide, 1991; Shen et al., 2021). Despite the inclusion of mesh optimization techniques (Hoppe et al., 1993) in modern mesh packages, this exhaustive method is still inclined to generate redundant complexities due to the discretization, needlessly burdening the computational load on a learning pipeline. Aiming for exact extraction, initial efforts focused on identifying linear regions (Serra et al., 2018) and employing Analytical Marching (Lei & Jia, 2020) to exactly reconstruct the zero-level isosurface. Contemporary approaches involve region subdivision, wherein the regions of the complex are progressively subdivided from neuron to neuron and layer to layer, allowing for the efficient calculation of the exponentially increasing linear regions (Raghu et al., 2017; Hanin & Rolnick, 2019; Humayun et al., 2023; Berzins, 2023). Nevertheless, many previous studies have explored CPWA functions consisting of fully connected layers and ReLU activation, primarily owing to the mathematical simplicity of hyperplanes defining linear regions. However, this restricts their applications in emerging domains of deep implicit surface learning that diverge from CPWA functions.
@@
The spectral bias of a multi-layer perceptron (MLP) hinders to effectively learn high frequencies, both theoretically and empirically (Tancik et al., 2020). To address this limitation, Fourier feature mapping, employing trigonometric projections similar to those used in the Transformer architecture (Vaswani et al., 2017), proves successful in representing complex 3D objects and scenes. Furthermore, to address its slow convergence, positional encoding methods and enhance rendering quality, novel positional encoding methods using trilinear interpolations, e.g., HashGrid (Müller et al., 2022), TensoRF (Chen et al., 2022) are introduced. Both techniques generate positional features through trilinear interpolation among the eight nearest corner features derived from a pre-defined 3D grid. These features are obtained by hashing from multi-resolution hash tables or factorized representations of a feature space, respectively. However, these approaches cause the function represented by neural networks to deviate from CPWA functions, as trilinear interpolation is not an affine transformation. Consequently, this results in the division of complex regions by intricate hypersurfaces.
@@
An eikonal  [^1] The eikonal is derived from the Greek, meaning image or icon. equation is a non-linear first-order partial differential equation governing the propagation of wavefronts. Notably, this equation finds application in the regularization of a signed distance function (SDF) in the context of neural surface modeling (Gropp et al., 2020; Wang et al., 2021; Li et al., 2023). If $\Omega$ represents a subset of the space ${\mathcal{X}}$ equipped with the metric $d$, the SDF $f({\bm{x}})$ is defined as follows:
@@
where $d$ represents the shortest distance from ${\bm{x}}$ to $\partial\Omega$ in the Euclidean space of ${\mathcal{X}}$. In the context of the Euclidean space with a piecewise smooth boundary, the SDF exhibits differentiability almost everywhere, and its gradient adheres to the eikonal equation. Specifically, for points ${\bm{x}}$ on the boundary, $\nabla f({\bm{x}})=N({\bm{x}})$ where the Jacobian of $f$ represents the outward normal vector, or equivalently, $|\nabla f({\bm{x}})|=1$.
@@ 3 Preliminaries
In this section, we present an overview of tropical geometry (Section 3.1) to provide a formal definition of the tropical hypersurface (Section 3.2). This hypersurface is significant as it serves as the decision boundary for ReLU neural networks optimizing an SDF, resulting in the formation of a polyhedral complex. For this, we delve into the examination of the tropical algebra of neural networks (Section 3.3). Furthermore, we discuss the edge subdivision algorithm (Berzins, 2023) for extracting the tropical hypersurface (Section 3.4), and we extend this discussion to include considerations for trilinear interpolation in Section 4. Given that Section 3.1 to 3.3 offer insights into tropical geometry, illustrating how each neuron contributes to forming the polyhedral complex, you may choose to skip these subsections if you are already acquainted with these concepts.
@@ 3.1 Tropical Geometry
Tropical  [^2] The unusual adjective tropical is named in honor of Hungarian-born Brazilian computer scientist Imre Simon (Katz, 2017). geometry (Itenberg et al., 2009; Maclagan & Sturmfels, 2021) describes polynomials and their geometric characteristics. As a skeletonized version of algebraic geometry, tropical geometry represents piecewise linear meshes using tropical semiring and tropical polynomial.
3.2 Tropical Hypersurface
@@
$V(f)=\{{\bm{x}}\in\mathbb{R}^{d}:c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}=c_{j}{\bm{x}}^{\mathrm{\alpha}_{j}}=f({\bm{x}})~{}~{}~{}~{}~{}$
$~{}~{}~{}~{}~{}\text{for some}~{}~{}~{}\mathrm{\alpha}_{i}\neq\mathrm{\alpha}_{j}\}$ (3)
where the tropical monomial $c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}$ with $d$-variate ${\bm{x}}$ is defined as $c\odot x_{1}^{a_{i,1}}\odot\cdots\odot x_{d}^{a_{i,d}}$ with $d$-variate $\mathrm{\alpha}_{i}=(a_{i,1},\dots,a_{i,d})$. In other words, this is the set of points where $f(x)$ equals two or more monomials in $f$.
3.3 Tropical Algebra of Neural Networks
@@
Although the neural networks $\nu$ is generally non-convex, it is known that the difference of two tropical signomials can represent $\nu$ (Zhang et al., 2018). A tropical signomial takes the same form as a tropical polynomial (ref. Definition 3.1), but it additionally allows the exponentials $\mathrm{\alpha}$ to be real values rather than integers. With tropical signomials $F$, $G$, and $H$,
$\nu^{(l)}({\bm{x}})$ $=H^{(l)}({\bm{x}})-G^{(l)}({\bm{x}})$ (5)
$\sigma^{(l)}\nu^{(l)}({\bm{x}})$ $=F^{(l)}({\bm{x}})-G^{(l)}({\bm{x}})$ (6)
where $F^{(l)}$, $G^{(l)}$, and $H^{(l)}$ are as follows:
$F^{(l)}({\bm{x}})$ $=\max\big{(}H^{(l)}({\bm{x}}),~{}G^{(l)}({\bm{x}})\big{)}$ (7)
@@
${\mathcal{B}}\subseteq\big{\{}{\bm{x}}:\nu_{j}^{(i)}({\bm{x}})=H_{j}^{(i)}({\bm{x}})-G_{j}^{(i)}({\bm{x}})=0,~{}~{}~{}$
$i\in(1,\dots,L),~{}j\in(1,\dots,H)\big{\}}$ (10)
where the subscription $j$ denotes the $j$-th element, assuming the hidden size of neural networks is $H$, while the $L$-th layer has a single output, for simplicity. Remind that tropical operations satisfy the usual laws of arithmetic for recursive consideration.
This implies that piecewise hyperplanes corresponding to the preactivations $\{\nu_{j}^{(i)}\}$ are the candidates shaping the decision boundary. Each linear region has the same pattern of zero masking by $\textrm{ReLU}(\cdot)$, effectively making it linear as an equation of a plane. In turn, if the pattern changes as an input is continually changed, it moves to another linear region. From a geometric point of view, this is why the number of linear regions grows polynomially with the hidden size $H$ and exponentially with the number of layers $L$ (Zhang et al., 2018; Serra et al., 2018).
@@ 3.4 Edge Subdivision for Tropical Geometry
Polyhedral  [^3] In geometry, a polyhedron is a three-dimensional shape with polygonal faces, having straight edges and sharp vertices. complex derivation from a multitude of linear regions is inefficient and may be infeasible in some cases. Rather, Berzins (2023) argue that tracking vertices and edges of the decision boundary sequentially considering the polynomial number of hyperplanes is particularly efficient, even enabling parallel tensor computations.
Initially, we start with a unit cube of eight vertices and a dozen edges, denoted by ${\mathcal{V}}$ and ${\mathcal{E}}$, respectively. For each piecewise linear, or folded in their term, hyperplane, we find the intersection points of each one of the edges and the hyperplane (if any) to add to ${\mathcal{V}}$. The divided edges by the intersection are also added to ${\mathcal{E}}$. Note that we find new edges of polygons, the intersections of convex polyhedra representing corresponding linear regions (Grigsby & Lindsey, 2022), and the folded hyperplane. Repeating these for all folded hyperplanes, before selecting all vertices ${\bm{x}}^{\star}$ where $\nu({\bm{x}}^{\star})=0$. Then, the corresponding edges would represent the decision boundary.
To elaborate with details, the order of subdivising is invariant as stated in Proposition 3.3; however, we must perform the subdivision with all hyperplanes in the previous layers in advance if we want to keep the current set of edges not crossing linear regions. One critical advantage of this principle is that we can find the intersection point by evaluating the output ratio of the two vertices of an edge, $\nu({\bm{x}}_{0})$ and $\nu({\bm{x}}_{1})$, which are proportional to the distances to the hyperplane by Thales’s theorem (Friedrich et al., 2008). The intersection point would simply be $\hat{{\bm{x}}}_{0,1}=(1-w){\bm{x}}_{0}+w{\bm{x}}_{1}$ where $w=|d_{0}|/|d_{0}-d_{1}|$ and $d_{k}=\nu_{j}^{(l)}({\bm{x}}_{k})$, upon the fact that $d_{0}\cdot d_{1}<0$ if the hyperplane divides the edge.
@@
They argue that the edge subdivision provides the optimal time and memory complexities of ${\mathcal{O}}(|{\mathcal{V}}|)$, linear in the number of vertices, while it can be efficiently computed in parallel. According to Hanin & Rolnick (2019), the number of linear regions is known for $o(N^{L}/L!)$ where $N$ is the total number of neurons, or $\phi(L,1)$ in our notation.
@@ 4.1 Piecewise Trilinear Networks
Let a unit hypercube be in the $D$-dimensional space, where its $2^{D}$ corners are represented by ${\bm{H}}\in\mathbb{R}^{2^{D}\times F}$. Given a weight ${\bm{w}}\in[0,1]^{D}$, the interpolation $\psi$ is defined as:
@@
where the interpolating weight $\omega(i,{\bm{w}})\in\mathbb{R}$ is defined using a left-aligned and zero-trailing binarization function $\textrm{B}\in\{0,1\}^{D}$ as follows:
@@
Notice that trilinear interpolation is linear for each dimension to prove it. The detailed proof can be found in Lemma A.1. ∎
@@
Let a positional encoding module be $\tau:\mathbb{R}^{D}\rightarrow\mathbb{R}^{F}$ using the trilinear interpolation of spatially nearby learnable vectors on a three-dimensional grid, e.g., HashGrid (Müller et al., 2022) or TensoRF (Chen et al., 2022). Usually, they transform the input coordinate ${\bm{x}}$ to ${\bm{w}}$ with a multi-resolution scale and find a relative position within a grid cell. Then, we define trilinear neural networks $\tilde{\nu}$, by prepending $\tau$ to the neural networks $\nu$ from Definition 3.2. The input and output dimensions of $\nu$ are adapted to $F$ and $1$, respectively.
@@
Note that $\tau$ is grid-wise trilinear. Using Lemma 4.2, $\tilde{\nu}$ is trilinear within the composite intersections of linear regions of $\nu$ and trilinear regions of $\tau$ if available. (We will discuss how to access the cubic corner features in the case that trilinear regions are no longer cubic in Section 5.3.) Then, we regard $\tilde{\nu}$ as piecewise trilinear.

### v2

@@ 1 Introduction
Recent advancements in visualizing deep neural networks [1, 2, 3, 4] significantly contribute to understanding their intricate structures. This progress provides valuable insights into the expressivity, robustness, training methodologies, and distinctive geometry of neural networks. By leveraging the inherent piecewise linearity in the Continuous Piecewise Affine (CPWA) functions, e.g., ReLU neural networks, each region of the input space is represented as a convex polyhedral. The assembly of these sets constructs a polyhedral complex that delineates the decision boundaries of neural networks [5].
@@
Building upon these breakthroughs, we explore the exciting prospect of analytically extracting a mesh representation from neural implicit surface networks [6, 7, 8]. The extraction of a mesh not only offers a precise visualization and characterization of the network’s geometry but also does so efficiently through the utilization of vertices and faces, in contrast to sampling-based methods [9, 10]. The sampling-based methods are limited by their black-box nature, which can obscure the underlying structure and lead to inefficiencies in capturing fine geometric details.
@@
These networks typically learn a signed distance function (SDF) by utilizing ReLU neural networks. However, recent approaches incorporate non-linear positional encoding techniques, such as trigonometric functions [11] or trilinear interpolations using hashing [12] or tensor factorization [13], to mitigate issues like spectral bias [14], ensuring fast convergence, and maintaining high-fidelity. Consequently, applying mesh extraction techniques based on CPWA functions becomes challenging.
Figure 1: The mesh is analytically extracted from piecewise trilinear networks, comprising both HashGrid [12] and ReLU neural networks, that have been trained to learn a signed distance function with the eikonal loss. We start with initial vertices and edges defined using the grid marks (1), and its linear subdivision (2a); however, intersecting polygons (ref. Section 3.2) need bilinear subdivision (2b) and consequentially trilinear subdivision (2c). Note that the linear and bilinear subdivisions are specific instances of trilinear subdivisions with straightforward solutions.
@@
Focusing on the successful and widely-adopted trilinear interpolating methods, we present novel theoretical insights and a practical methodology for precise mesh extraction. Following a novel review of the edge subdivision method [4], an efficient mesh extraction technique for CPWA functions, within the framework of tropical geometry (Section 3), we demonstrate that ReLU neural networks incorporating the trilinear module are piecewise trilinear networks (Definition 4.1). Then, we establish that the hypersurface within the trilinear region becomes a plane under the eikonal [15] constraint (Theorem 4.5), a common practice during the training of SDF.
@@
Building upon this observation, we propose a method for approximating the determination of intersecting points among three hypersurfaces. Given the inherent curvature of these surfaces, achieving an exact solution is infeasible. In our approach, we utilize two hypersurfaces and a diagonal plane for a more feasible and precise approximation (Theorem 4.7). Additionally, we introduce a methodology for arranging vertices composed of faces to implicitly represent normal vectors. We experimentally confirm accuracy and simplicity using chamfer distance, efficiency, and angular distance, simultaneously exploring the relationship between the eikonal loss and the planarity of the hypersurface. In Figure 1, we present the analytically generated skeleton and normal map produced by our method using piecewise trilinear networks composed of HashGrid and ReLU neural networks. For a more detailed view, please refer to Figure 5 in Appendix.
@@
    We present novel theoretical insights and a practical methodology for precise mesh extraction, from the piecewise trilinear networks using the HashGrid.
@@
    We validate its correctness and parsimony through chamfer distance and efficiency, and angular distance, showing the correlation between the eikonal loss and hypersurface planarity.
@@ 2 Related work
A conventional black-box approach is to assess a given function at sampled points to derive the complex of the function, commonly through marching cubes [9] or marching tetrahedra [16, 10]. Despite the inclusion of mesh optimization techniques [17] in modern mesh packages, this exhaustive method is still inclined to generate redundant complexities due to the discretization, needlessly burdening the computational load on a learning pipeline. Aiming for exact extraction, initial efforts focused on identifying linear regions [18] and employing Analytical Marching [2] to exactly reconstruct the zero-level isosurface. Contemporary approaches involve region subdivision, wherein the regions of the complex are progressively subdivided from neuron to neuron and layer to layer, allowing for the efficient calculation of the exponentially increasing linear regions [19, 20, 21, 4]. Nevertheless, many previous studies have explored CPWA functions consisting of fully connected layers and ReLU activation, primarily owing to the mathematical simplicity of hyperplanes defining linear regions. However, this restricts their applications in emerging domains of deep implicit surface learning that diverge from CPWA functions.
@@
The spectral bias of a multi-layer perceptron (MLP) hinders effectively learning high frequencies, both theoretically and empirically [14]. To address this limitation, Fourier feature mapping, employing trigonometric projections similar to those used in the Transformer architecture [22], proves successful in representing complex 3D objects and scenes. Furthermore, to address its slow convergence, positional encoding methods and enhance rendering quality, novel positional encoding methods using trilinear interpolations, e.g., HashGrid [12], TensoRF [13] are introduced. Both techniques generate positional features through trilinear interpolation among the eight nearest corner features derived from a pre-defined 3D grid. These features are obtained by hashing from multi-resolution hash tables or factorized representations of a feature space, respectively. However, these approaches cause the function represented by neural networks to deviate from CPWA functions, as trilinear interpolation is not an affine transformation. Consequently, the complex regions are divided by intricate hypersurfaces.
@@
An eikonal  [^1] The eikonal is derived from the Greek, meaning image or icon. equation is a non-linear first-order partial differential equation governing the propagation of wavefronts. Notably, this equation finds application in the regularization of a signed distance function (SDF) in the context of neural surface modeling [23, 6, 8]. If $\Omega$ represents a subset of the space ${\mathcal{X}}$ equipped with the metric $d$, the SDF $f({\bm{x}})$ is defined as follows:
@@
where $d$ represents the shortest distance from ${\bm{x}}$ to $\partial\Omega$ in the Euclidean space of ${\mathcal{X}}$. In the Euclidean space with a piecewise smooth boundary, the SDF exhibits differentiability almost everywhere, and its gradient adheres to the eikonal equation. Specifically, for points ${\bm{x}}$ on the boundary, $\nabla f({\bm{x}})=N({\bm{x}})$ where the Jacobian of $f$ represents the outward normal vector, or equivalently, $|\nabla f({\bm{x}})|=1$.
@@ 3.1 Tropical geometry
Tropical  [^2] The tropical is named in honor of Hungarian-born Brazilian computer scientist Imre Simon [24]. geometry [25, 26] describes polynomials and their geometric characteristics. As a skeletonized version of algebraic geometry, tropical geometry represents piecewise linear meshes using tropical semiring and tropical polynomial (ref., Section A.1). The set of points where a tropical polynomial $f$ is non-differentiable, due to tropical sum $\oplus:=\max(\cdot)$, is called tropical hypersurface:
@@
$V(f)=\{{\bm{x}}\in\mathbb{R}^{d}:c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}=c_{j}{\bm{x}}^{\mathrm{\alpha}_{j}}=f({\bm{x}})~{}~{}~{}~{}~{}\text{for some}~{}~{}~{}\mathrm{\alpha}_{i}\neq\mathrm{\alpha}_{j}\}$
where the tropical monomial $c_{i}{\bm{x}}^{\mathrm{\alpha}_{i}}$ with $d$-variate ${\bm{x}}$ is defined as $c\odot x_{1}^{a_{i,1}}\odot\cdots\odot x_{d}^{a_{i,d}}$ with $d$-variate $\mathrm{\alpha}_{i}=(a_{i,1},\dots,a_{i,d})$ and tropical product :=$\odot$:=+⊙ := +.
@@
Although the neural networks $\nu$ is generally non-convex, it is known that the difference of two tropical signomials can represent $\nu$ [27] (ref., Section A.3). Given this observation, we proceed to examine the following proposition:
@@
${\mathcal{B}}\subseteq\big{\{}{\bm{x}}:\nu_{j}^{(i)}({\bm{x}})=0,~{}i\in(1,\dots,L),~{}j\in(1,\dots,H)\big{\}}$
where the subscription $j$ denotes the $j$-th element, assuming the hidden size of neural networks is $H$, while the final $L$-th layer has a single output for rather discussions as an SDF. Remind that tropical operations satisfy the usual laws of arithmetic for recursive consideration.
Given that piecewise hyperplanes corresponding to the preactivations, $\nu_{j}^{(i)}$ are the candidates shaping the decision boundary, it enables for-loop iterations over the neurons across layers in Algorithm 2, not visiting every exponentially growing linear region [27, 18]. We note that this formal explanation through tropical geometry is unprecedented in the previous work [4].
@@ 3.2 Edge subdivision for tropical geometry
Polyhedral  [^3] In geometry, a polyhedron is a three-dimensional shape with polygonal faces, having straight edges and sharp vertices. complex derivation from a multitude of linear regions is inefficient and may be infeasible in some cases. Rather, Berzins [4] argue that tracking vertices and edges of the decision boundary sequentially considering the polynomial number of hyperplanes is particularly efficient, even enabling parallel tensor computations for edge subdivisions.
Initially, we start with a unit cube of eight vertices and a dozen edges, denoted by ${\mathcal{V}}$ and ${\mathcal{E}}$, respectively. For each piecewise linear, or folded in their term, hyperplane, we find the intersection points of each one of the edges and the hyperplane (if any) to add to ${\mathcal{V}}$. The divided edges by the intersection are also added to ${\mathcal{E}}$. Note that we also find new edges of polygons, the intersections of convex polyhedra representing corresponding linear regions [5], and the folded hyperplane. Repeating these for all folded hyperplanes, before selecting all vertices ${\bm{x}}^{\star}$ where $\nu({\bm{x}}^{\star})=0$. Then, the valid edges connecting those vertices would represent the decision boundary.
To elaborate with details, the order of subdividing is invariant as stated in Proposition 3.3; however, we must perform the subdivision with all hyperplanes in the previous layers in advance if we want to keep the current set of edges not crossing linear regions. One critical advantage of this principle is that we can find the intersection point by evaluating the output ratio of the two vertices of an edge, $\nu({\bm{x}}_{0})$ and $\nu({\bm{x}}_{1})$, which are proportional to the distances to the hyperplane by Thales’s theorem [28]. The intersection point would simply be $\hat{{\bm{x}}}_{0,1}=(1-w){\bm{x}}_{0}+w{\bm{x}}_{1}$ where $w=|d_{0}|/|d_{0}-d_{1}|$ and $d_{k}=\nu_{j}^{(l)}({\bm{x}}_{k})$, upon the fact that $d_{0}\cdot d_{1}<0$ if the hyperplane divides the edge.
@@
Berzins [4] argue that the edge subdivision provides the optimal time and memory complexities of ${\mathcal{O}}(|{\mathcal{V}}|)$, linear in the number of vertices, while it can be efficiently computed in parallel. According to Hanin and Rolnick [20], the number of linear regions is known for $o(N^{L}/L!)$ where $N$ is the total number of neurons, or $\phi(L,1)$ in our notation.
@@ 4.1 Piecewise trilinear networks
Let a unit cube be in the $D$-dimensional space, where its corners are represented by ${\bm{H}}\in\mathbb{R}^{2^{D}\times F}$. Given weights ${\bm{w}}\in[0,1]^{D}$, the $\psi$ is defined as:
@@
where the interpolating weight $\omega(i,{\bm{w}})\in\mathbb{R}$ is defined using a left-aligned and zero-trailing binarization function $\textrm{B}\in\{0,1\}^{D}$ (ref., Section B.3) as follows:
@@
Notice that trilinear interpolation is linear for each dimension to prove it. The detailed proof can be found in Lemma D.1. ∎
@@
Let a positional encoding module be $\tau:\mathbb{R}^{D}\rightarrow\mathbb{R}^{F}$ using the trilinear interpolation of spatially nearby learnable vectors on a three-dimensional grid, e.g., HashGrid [12] or TensoRF [13]. Usually, they transform the input coordinate ${\bm{x}}$ to ${\bm{w}}$ with a multi-resolution scale and find a relative position within a grid cell. Then, we define trilinear neural networks $\tilde{\nu}$, by prepending $\tau$ to the neural networks $\nu$ from Definition 3.2. Here, the input and output dimensions of $\nu^{(L)}$ are $F$ and $1$ (an SDF distance), respectively.
@@
Note that $\tau$ is grid-wise trilinear. Using Lemma 4.2, $\tilde{\nu}$ is trilinear within the composite intersections of linear regions of $\nu$ and trilinear regions of $\tau$ if available. (We will discuss how to access the cubic corner features in the case that trilinear regions are no longer cubic in Section 5.) Then, we regard $\tilde{\nu}$ as piecewise trilinear.